OpenStudy (anonymous):

What is the slope of the line that passes through the points (-6,1) and (4,-4)? a) -2 b) 2 c) -1/2 d) 1/2

OpenStudy (anonymous):

rise over run, (-4-1)/(4+6) =1/2

OpenStudy (anonymous):

oops missing a minus sign.

OpenStudy (anonymous):

sorry, -1/2

OpenStudy (anonymous):

do this in your head: from -6 to 4 is ten units right. form 1 tp -4 is 5 units down. \[\frac{-5}{10}=-\frac{1}{2}\]

OpenStudy (anonymous):

oh no its -1/2 yeah lol